# Cocoa: Co-Planning and Co-Execution with AI Agents

### v3

@@ 6.1.2. Co-execution afforded steerability
The ability to directly edit the agent’s outputs via direct manipulation by deleting irrelevant items and adding items of their own also improved steerability. Many participants, including P6, liked the precision with which they could edit the output of a particular step, which is much more difficult in a chat interface: “With a conversation agent, even if I’m really good at prompting it, I have to redo the prompt and keep changing that original prompt.” (P6). This feature also allowed P2 to quickly pivot and repurpose the output when the agent produced unexpected (albeit helpful) paper search queries: “I was looking for more application intervention based stuff, but this is great because now I can actually use these keywords to broaden my horizon of how I was thinking about these solutions, which is I think interesting. I’m just trying to remove a couple [for future steps].” P15 thought direct manipulation enabled more granular control than chat, allowing them to easily guide the agent with their expertise: “[Cocoa ] was easier to control because [the outputs] are so specific. My control is scoped down to these very small tasks. If I want to add more papers, that’s easy to do.”
6.2. Task-Specific Preferences for Interactive Planning vs. Chat (L2)
It is reasonable for participants not to strictly prefer Cocoa over our chat baseline for every research task. After all, chat is a strong baseline because it is flexible, easy to use, and ubiquitous. Here, we aim to surface participants’ insights on when they prefer interactive plans over chat, and vice versa. Ultimately, many participants recognized that both interfaces have their strengths and weaknesses, and, in P13’s words, it was “not fair to say one is better than the other.” Instead, participants imagined using the two systems at different stages of their research. In Section 8.1, we further discuss the costs and benefits between the two paradigms and how to combine them in future systems to get the best of both worlds.
6.2.1. Preferences for interactive plans
Participants generally agreed that when tackling questions or tasks that required a structured or organized approach, they preferred Cocoa over chat. One example of such tasks is literature review and synthesis. P12 considered the plan more usable as an end product than a conversation: “I think it’s a better way to have a finished draft [plan] that I can directly use as opposed to trying to dilute a plan in a conversation.” Others appreciated that the plan helped them stay organized and consolidated lot of information: “[the plan] is a really good way to organize my thoughts” (P14) and “everything can be part of one consolidated [artifact] which I can refer to again” (P7).
The structure of an interactive plan was also helpful when participants had an idea of their desired final output, but were unsure of how to get there. For their task, P6 appreciated Cocoa because they needed “some concrete steps even if I need to iterate over those steps,” and because the agent’s first step “helped me get to search terms [for papers] which was what I was struggling with.” Similarly, P14 said that plans can help them overcome mental blockers when faced with a tough problem: “If I’m really stuck and frustrated, I can click a single button, and it could output a whole plan for me. That would make me feel really safe because it’s always here outputting something for me which potentially get me out of this frustration.” For P16, they did not consider anything but a structured approach to be acceptable for research. They envisioned the research process as “a lot of subtasks that you just come up with” and collected information to be “concrete items” that they would operate on with “high-level actions [like] directly extracting insights.” As a result, Cocoa aligned well with their mental model and found the chat “very frustrating […] and almost like the opposite” of their preferred approach.
6.2.2. Preferences for chat-based interactions
While interactive plans afford more steerability, chat-based interactions may be more appropriate when the task at hand is open-ended and difficult to create plans in advance, or too narrowly scoped to benefit from a multi-step plan. Many preferred chat for freeform exploration, where the nature of the desired final output was ambiguous. P2, while conducting an early brainstorming task, thought it was “a little overwhelming to go through all these steps, and I also felt this kind of pressure to stick with the plan although I was editing it.” P1 suggested that, when brainstorming, Cocoa could show them “a single step at a time” and allow them to “choose from options for the next step” after they have seen the outputs for the previous step. Question-answering was another category of tasks where chat was preferred, where the next step depends on the output of the previous step. P13 found it much easier to “keep on asking [follow-up] questions” to drill down on a specific line of inquiry in chat.
6.3. Task Assignment to Agent vs. User (L3)
In our formative study (Section 3.2.3), participants shared that they preferred to assign tasks related to information retrieval and brainstorming tasks to the agent while keeping higher-level reasoning and synthesis tasks for themselves. In our study, however, most participants toggle all plan steps—even higher-level ones assigned to them by default—to the agent. Some participants chose not to include user steps in their plans because they were curious about the agent’s capabilities and limitations—a natural tendency when learning to use a new system. For example, P1 wanted to see “how the bot does it” while acknowledging that they prefer to “make the decisions and here, these are more exploratory steps rather than decision-making ones.” P13 also opted for this approach and considered the agent’s errors to be low-stakes because they can directly modify the outputs at any time: “when I don’t like something, I can just easily remove it.” P9 also felt assured that they could edit the outputs and did not consider the output quality to be particularly problematic: “usually as long as it’s somewhat helpful and it’s just running in the background, I would take a look at [the output].” Adopting a cost-benefit framework, they mentioned that they would leave a step as an agent step as long as “the cost of looking at the agent’s output is less than the benefit of whatever insights I obtain.”
While we did not seen substantial use of user steps in the lab, participants also did not view the ability to create user steps negatively. Many speculated in the interviews that they would leverage this mechanism more if they had more time to interact with the system and outside of the lab environment. P8 explicitly said that “I assign everything to [the agent] because I don’t think we have that much time for me to refine my [own steps].” This further motivated us to conduct a longitudinal deployment study of Cocoa, and Section 7.2.1 details how participants’ task assignment strategy differs in the field.
@@ 7. Deployment Study and Results
As surfaced in Section 6.3, participants’ usage behaviors in Cocoa may shift given more time to interact with the system. To obtain a more holistic understanding of how Cocoa can support researchers in their day-to-day work, we conducted a 7-day field deployment study with 7 participants.
@@ 7.1. Participants and Procedure
Participants used Cocoa to work on a real-world, in-progress research project of their choice over a 7-day period. Before the study, the study facilitator sent each participant a user manual documenting Cocoa’s functionalities. The manual also contained some general guidelines for the week (see Appendix J), including aiming to spend at least 90 minutes with the system throughout the duration of the study. Each participant could log onto their personal workspace in Cocoa with a set of credentials provided by the study facilitator. We used system logs to verify that there was at least semi-regular activity for each participant throughout the study.
@@ 7.2. Field Deployment Study Results
7.2.1. Step assignment strategy changed: more user steps for complex longitudinal tasks
Compared to the lab study, where most co-execution was focused on removing or adding to an agent step’s output, participants in the field study assigned more plan steps to themselves. These user steps often involve high-level strategic decision-making, writing and developing arguments, running experiments and evaluations, deep reading of literature, and articulating core research ideas and novelty. The ability to assign steps to the agent and themselves helped participants create a back-and-forth workflow with the agent. For example, P15 first figured out “broad buckets of literature or themes or topics that I need to talk about” before passing those to the agent to “help me brainstorm ways to connect those bodies of literature.” P3, who had already conducted some interviews for their project, first had the agent look up key literature before assigning themselves a step to “pass in more of the dynamics from my interview notes” and then had the agent propose interventions based on the literature and notes. For many participants, user steps came later in the plan because of their high-level nature. P11 often assigned the last plan step to themselves because “it was about the main innovation in the project idea […] no one has [done it] before so I don’t trust the agent.” P16 assigned themselves the last step as a to-do item and a reminder to reason about how the agent’s output “could actually be tied into some of the document.” Participants also indicated that their step assignment also depended on the stage of research they were in. P3 admitted that “if I was just doing initial discovery and study planning, I would have gone more agent [steps].”
7.2.2. Interleaving co-planning and co-execution was valuable
Participants identified the interleaving of co-planning and co-execution afforded by interactive plans as a unique and valuable aspect of Cocoa. This interleaving allowed for a flexible and iterative workflow. P12 shared that “I don’t need to run the [earlier] retrieval step again. I know that those papers are relevant” but instead could focus on “only improving the second step […] to identify something different from the literature.” Similarly, P16 felt like they could easily “run the plan fully and then go in and adjust the rerun things,” which enabled them to “debug” suboptimal parts of the plan: “by forcing these checkpoints of brainstorming search queries and assembling lists of papers, it helps me get a sense of what parts of the process are good and which parts need work.” Outputs from co-execution also inspired new approaches for participants to “iteratively build upon a plan or reshape a plan” (P5). Participants even mentioned that they wanted similar interleaving strategies integrated into other LLM-powered systems; for P11, it was “one of the improvements I want to see in LLM [reasoning] models” so that users can force an early stop to the reasoning or steer it in more productive directions.
@@ 7.2.3. Cocoa was most useful for literature synthesis and early-stage project planning
All participants agreed that Cocoa was especially helpful for literature discovery and synthesis. P15 found the plan steps involving literature search “super useful” and the search query generation steps that commonly preceded them to be “really helpful for refining my thought process because it made me think more carefully about what I’m actually trying to look for.” P10 also found Cocoa to be “really helpful with looking for papers” and that it was “less burdensome” to use compared to Google Scholar. The combination of plan steps involving search query generation and paper search using those queries left P10 with “more faith that I was getting a good coverage of the field.” P3 was impressed by Cocoa’s performance on thematic summarization steps: “[Cocoa] did a really good job at grouping papers that went together.” P11 also appreciated Cocoa’s speed at literature search and synthesis and found it conducive to early-stage explorations: “if I were to come up with a summary of 20 papers, it takes probably a day, but [in Cocoa] it’s around a minute which is really useful to formulate early-stage research ideas.” Beyond literature search, the structure of the plan was also useful to “break down tasks and provide a sense of progress” (P16) and “identifying important actionable steps I could take with a research problem” (P12). Overall, participants found Cocoa valuable in their work, with P3 in particular sharing that “I have an entire tab in my Google doc where we have stuff from Cocoa.”
@@ 7.2.4. Additional desiderata
Participants identified several desiderata for Cocoa after spending more time with it. Some participants wanted to clearly see what context Cocoa was used to generate plans and execute plan steps. For example, P10 was unsure whether the system was “using the context from my other plans to generate the [other] plan.” The system indeed was, but did not visualize relevant context in the document. P3, P15, and P16 all wanted more transparency into the shared context pool (Section 4.2) used across all steps and the specific context drawn upon by each step. Several participants desired tool interoperability, such that they could swap out Cocoa’s default tools for their own ones (e.g., an LLM finetuned on paper summaries instead of Cocoa’s paper summarizer). This way, interactive plans would serve as a “hub” for their custom tools. Finally, P5 took inspiration from Notion’s ability to nest pages to suggest nested plans, where plans can be organized and expanded in a hierarchical manner. These suggestions all provide valuable directions for future work.
@@ 8. Discussion
8.1. Chat, Interactive Plans, or Both?
In practice, rather than choosing between either interactive plans or chat when building future systems, we see research opportunities for AI agent interfaces to strategically combine the two, especially when assisting the user with complex, long-running requests.
@@ 8.1. Chat, Interactive Plans, or Both?
Consider “deep research” systems that assemble a detailed report on a particular topic upon searching for and synthesizing information from around the web (Google, 2025; OpenAI, 2025a; Perplexity, 2025). Some of these systems may clarify user requests in a chat conversation (OpenAI, 2025a) or allow the user to prompt the agent to devise a different plan (Google, 2025) before starting the research process, but neglect user involvement thereafter. Integrating interactive plans can be particularly valuable here because they afford iteration and control over specific parts of planning and execution. For example, the user may not be satisfied with the sources the agent is using to inform its answers and wants to steer the agent to use information from a narrow set of websites for a particular plan step. This kind of interaction is expensive without interactive plans because it requires re-running the entire research process. Interactive plans allow the user to focus on improving outcomes of a particular step by editing that step and rerunning it or directly editing the agent’s outputs at that step. Rather than re-executing the entire plan, the agent updates only the subsequent plan steps with the new user input. After the interactive plan has finished executing, users may return to interacting with the agent in a chat interface to ask questions about specific outputs or steps. These are precisely the benefits that computational notebooks provide for data analysts (Choi et al., 2023; Ayobi et al., 2023; Chattopadhyay et al., 2020).
In addition to using chat and interactive plans sequentially, we can also consider how they can be integrated within one another. Embedding chat within interactive plans can provide more flexibility and organization. For example, during co-planning, chat-based features can be used to specify higher-level desiderata for the proposed plans when none are satisfactory for the user. During co-execution, the user may engage in conversation with the agent at every step to iterate on outputs. This way, the interactive plan also serves as a way to organize conversation threads with the agent as the agent operates over many plan steps. Conversely, interactive plans can be incorporated into chat-first interfaces to provide more transparency and steerability. This can be useful for open-ended exploratory problems, with occasional periods of structured execution.
8.2. The Practical Significance of User Steps and Advanced Planning
@@
In Cocoa, users are presented with a multi-step interactive plan up front, which the agent adheres to and can modify while operating. This plan also allows users to delegate a step to themselves or the agent. Both design decisions are a departure from many existing agentic AI systems, where the agent dynamically constructs a plan step-by-step as it generates outputs (Yao et al., 2022b) and seeks user input only when it is unable to complete a task (Wu, 2024). We discuss two practical reasons—cost and oversight—for why such a departure may be desirable for practical agents in the real world.
@@
AI agents can be prohibitively expensive to run (Kapoor et al., 2024; Yang et al., 2024). Because an agent’s actions often involve multiple LLM calls, costs can quickly accumulate, especially for more complex and long-running tasks. In response, academic projects have implemented cost-capping measures to alleviate this (Yang et al., 2024), and researchers have called for agent evaluations to be cost-controlled (Kapoor et al., 2024). In user-facing interactive systems, cost also includes time spent waiting for agents to come back with results. These cost considerations change user behavior. When participants made the decision to assign most or all steps to the agent in our lab studies, they viewed the agent’s actions as rather inconsequential and low-cost—they could run the agent and see what happens, and edit the output post hoc (Section 6.3). However, while this was the case in the lab, it may not be true in some real-world contexts where users and/or developers need to pay monetary and time costs for each agent step. The user steps become an important cost-saving measure. User can choose to complete a step that is not particularly burdensome or is too difficult for current AI agents to complete effectively. Efficiently leveraging human effort and expertise where appropriate via user steps can be a key step towards developing cost-aware agents that are usable and affordable in practice.
@@
The increased autonomy of AI agents also comes with increased risks, as AI harms become more difficult to anticipate, and accountability for AI actions becomes harder to trace (Chan et al., 2023, 2024). Interactive plans not only provide more transparency into agent actions, but also provide some assurance about and control over the agent’s execution trajectory. By generating a plan ahead of execution that the agent adheres to, harm anticipation and mitigation become more tractable. While step-by-step plan generation used by existing agent frameworks, such as ReAct (Yao et al., 2022b), may also accomplish a similar goal by asking for user approval whenever a new step is generated, this can severely limit agent autonomy when desirable and may also disengage the user if their only form of supervision is repetitive rubber-stamping (Kapoor et al., 2024). Additionally, advanced planning enables assessments of task risk pre-execution. Plan steps determined to be of higher risk (e.g., requiring working with passwords) can be assigned to the user—even if the agent is capable of completing them—as a risk mitigation measure.

### v4

@@ 6.2.2. Fluid movement between co-planning and co-execution enabled granular steering
Interleaving co-planning and co-execution via direct manipulation in particular also contributed to granular steering. Many participants, including P6, liked the precision with which they could edit the output of a particular step compared to a chat interface: “With a conversation agent, even if I’m really good at prompting it, I have to redo the prompt and keep changing that original prompt.” (P6). This feature also allowed P2 to quickly re-plan and repurpose the output when the agent produced unexpected (albeit helpful) paper search queries: “I was looking for more application intervention based stuff, but this is great because now I can actually use these keywords to broaden my horizon of how I was thinking about these solutions, which is I think interesting. I’m just trying to remove a couple [for future steps].” Finally, participants found that interleaved co-planning and co-execution scoped tasks in a way that naturally afforded more precise control, especially when paired with direct manipulation, as P15 shared: “[Cocoa ] was easier to control because [the outputs] are so specific. My control is scoped down to these very small tasks. If I want to add more papers, that’s easy to do.”
@@ 7. Deployment Study and Results
In addition to interleaving co-execution and co-planning, we were also interested in studying the delegation of human and AI work. Thus, we tracked which steps participants assigned to the agent vs. themselves.
In our formative study (Section 3.2.3), participants preferred to assign tasks related to information retrieval and brainstorming to the agent while keeping higher-level reasoning and synthesis tasks for themselves. In our lab study, however, most participants toggled all plan steps—even higher-level ones assigned to them by default—to the agent. Some avoided user steps because they were curious about “how the bot does it” (P1). Understandably, testing the agent’s capabilities and limitations is a natural tendency when learning to use a new system. Many speculated that they would leverage user steps if they had more time to interact with the system outside of the lab environment. P8 explicitly said that “I assign everything to [the agent] because I don’t think we have that much time for me to refine my [own steps].”
To obtain a more holistic understanding of how Cocoa can facilitate flexible delegation of human-AI work and more generally assist researchers day-to-day, we conducted a 7-day field deployment study with 7 participants.
@@ 7.1. Participants and Procedure
Participants used Cocoa to work on a real-world, in-progress research project of their choice over a 7-day period. Before the study, the study facilitator sent each participant a user manual [^12] See Supplementary Materials. documenting Cocoa’s functionalities. The manual also contained some general guidelines for the week (see Appendix I), including aiming to spend at least 90 minutes with the system throughout the duration of the study. Each participant could log onto their personal workspace in Cocoa with a set of credentials provided by the study facilitator. We used system logs to verify that there was at least semi-regular activity for each participant throughout the study.
@@ 7.2.1. Cocoa was most useful for literature synthesis and early-stage project planning
All participants agreed that Cocoa was especially helpful for literature discovery and synthesis. This illustrates the value of interleaving human-AI efforts across planning and execution in complex information foraging and sensemaking activities. P15 found the structured plan for conducting literature search to be “really helpful for refining my thought process because it made me think more carefully about what I’m actually trying to look for.” P10 agreed, and also felt that the combination of plan steps involving search query generation and paper search left them with “more faith that I was getting a good coverage of the field.” P11 also appreciated Cocoa’s speed at literature search and synthesis and found it conducive to early-stage explorations: “if I were to come up with a summary of 20 papers, it takes probably a day, but [in Cocoa] it’s around a minute which is really useful to formulate early-stage research ideas.” Beyond literature search, the structure of the plan was also useful to “break down tasks and provide a sense of progress” (P16) and “identifying important actionable steps I could take with a research problem” (P12). Overall, participants found Cocoa valuable in their work, with P3 in particular sharing that “I have an entire tab in my Google doc where we have stuff from Cocoa.”
7.2.2. Step assignment strategy changed: more delegation of work with longitudinal use
Compared to the lab study, where steps were mostly assigned to the agent, participants in the field study assigned more plan steps to themselves. This was likely due to a combination of increased familiarity with the agent’s capabilities and system features, as well as more opportunity to reason about which steps they wanted to keep for themselves and why. These user steps often involved high-level strategic decision-making, writing and developing arguments, running experiments and evaluations, deep reading of literature, and articulating core research ideas and novelty.
The ability to assign steps to the agent and themselves helped participants create a back-and-forth workflow with the agent. For example, P15 first figured out “broad buckets of literature or themes or topics that I need to talk about” before passing those to the agent to “help me brainstorm ways to connect those bodies of literature.” P3, who had already conducted some interviews for their project, first had the agent look up key literature before assigning themselves a step to “pass in more of the dynamics from my interview notes” and then had the agent propose interventions based on the literature and notes. For some participants, user steps came later in the plan because of their high-level nature. P11 often assigned the last plan step to themselves because “it was about the main innovation in the project idea […] no one has [done it] before so I don’t trust the agent.”
Participants also indicated that their step assignment also depended on the stage of research they were in. P3 admitted that “if I was just doing initial discovery and study planning, I would have gone more agent [steps].” These findings suggest that participants benefitted from the ability to flexibly configure agency in their human-AI workflows.
Figure 10. Examples from our field deployment where user and agent steps were interleaved in participants’ plans. The plans steps were rewritten to redact project-specific details.
Two examples of plans in Cocoa featuring alternating user- and agent-assigned steps, suggesting that user and agent efforts were interleaved.
7.2.3. Interleaving co-planning and co-execution supported evolving research trajectories
Just like in the lab study, participants valued interleaving of co-planning and co-execution, but the value for supporting flexible and iterative workflow became more apparent with longitudinal use. As participants saw the results of execution, their understanding of the problem space shifted, which led them to want to anchor certain steps and outputs as “checkpoints,” and use those as a springboard for exploring new territory. For example, P12 shared that “I don’t need to run the [earlier] retrieval step again. I know that those papers are relevant” but instead could focus on “only improving the second step […] to identify something different from the literature.” Similarly, P16 felt like they could easily “debug” suboptimal parts of the plan: “by forcing these checkpoints of brainstorming search queries and assembling lists of papers, it helps me get a sense of what parts of the process are good and which parts need work.” Outputs from co-execution also inspired new approaches for participants to “iteratively build upon a plan or reshape a plan” (P5). Participants even mentioned that they wanted similar interleaving strategies integrated into other LLM-powered systems; for P11, it was “one of the improvements I want to see in LLM [reasoning] models” so that users can force an early stop to the reasoning or steer it in more productive directions. Overall, our deployment of Cocoa suggests that interleaving co-planning and co-execution helps support the non-linear nature of scientific inquiry.
@@ 7.2.4. Additional desiderata
Participants identified several desiderata for Cocoa after spending more time with it. Some participants wanted to clearly see what context Cocoa used to generate plans and execute plan steps. For example, P10 was unsure whether the system was “using the context from my other plans to generate the [other] plan.” The system indeed was, but did not visualize relevant context in the document. P3, P15, and P16 all wanted more transparency into the shared context pool (Section 4.2) used across all steps and the specific context drawn upon by each step. Several participants desired tool interoperability, such that they could swap out Cocoa’s default tools for their own ones (e.g., an LLM finetuned on paper summaries instead of Cocoa’s paper summarizer). This way, Cocoa would serve as a “hub” for their custom tools. Finally, P5 took inspiration from Notion’s ability to nest pages to suggest nested plans, where plans can be organized and expanded in a hierarchical manner. These suggestions all provide valuable directions for future work.
@@ 8. Discussion
8.1. Design Considerations for Co-Agency in Human-Agent Collaboration
Researchers have called for reframing the design of human-AI systems as not merely the specification of affordances, but construction of open design spaces that support human-AI co-agency (Satyanarayan, 2024). That is, rather than providing a fixed configuration of human and AI agency across all interactions, users can flexibly vary the agency they wish assert and have the AI modulate its agency in response. Cocoa offers an early instantiation of this vision. We now reflect upon several key design decisions and discuss how future systems may build upon them or explore alternatives when designing for co-agency in human-agent systems.
We expect that fluid re-planning will continue to be important and challenging to design for in future systems. Plans are bound to shift as the user reviews agent outputs and becomes inspired with new ideas. For example, researchers in both our lab and deployment studies dynamically re-planned with the agent upon gaining more clarity on what they actually wanted to explore after co-executing earlier steps (see Sections 6.2 and 7.2.3). In Cocoa, re-planning replaces outdated steps with new ones. However, participants sometimes expressed a desire to preserve the old steps because their outputs can still be useful to reference in the future. We thus see plan branching as a promising design pattern for exploring counterfactual plan paths. For example, systems can allow users to swap certain steps in and out of the plan, or set checkpoints at specific steps to anchor exploration. By doing so, useful work can be better retained while supporting flexible plan redirection.
Relatedly, the length of the plan should be calibrated depending on the task. Some tasks, such as open-ended brainstorming, benefit from shorter, partial plans where next steps can be flexibly determined based on prior execution. Indeed, a few participants expressed a preference for a step-by-step planning approach over Cocoa’s default plans when conducting brainstorming, where they did not know what next step would be until they saw the results of the previous step. On the other hand, for tasks with more structured and predictable workflows, such as literature review, participants appreciated interacting with a full plan. We calibrated the lengths of plans generated by Cocoa based on participants’ plans in our formative study. We recommend designers of human-agent systems to consider what the appropriate plan length is for tasks they hope to support, and calibrate accordingly.
Finally, user expertise and upskilling potential should be considered when designing for an appropriate level of agent control. As surfaced in our formative study (Section 3.2.3), there are two primary reasons why users prefer keeping a task for themselves over delegating it to the agent. First, the user can perform the task more effectively than the agent. This is often the case where the user holds more expertise (e.g., brainstorming research questions in a user’s primary research area). Second, the task is intrinsically valuable for the user to do themselves, even if an agent is capable of automating it (e.g., reading seminal papers to develop research taste). If a human-agent system is designed for expert users who want to regularly hone their skills and intuition, providing many levels of control throughout agent planning and execution (as in Cocoa) may be appropriate. If the system is for navigating unfamiliar domains with agents and users are not looking to upskill, fewer controls may enhance user experience. Because interactions such as co-planning and co-execution can burden the user with additional cognitive load, determining the user’s capacity and willingness to take on more cognitive load can inform the appropriate level of control to afford.
These design considerations were drawn from the specific definition of agency (from Bennett et al. (Bennett et al., 2023)) and domain (scientific research) specific to our study. The design space is vast, and we are excited to witness the application, refinement, and even refutation of these considerations as future work navigates this space with alternative interpretations of agency in different domains.
8.2. The Agent Notebook Paradigm?
The design of Cocoa’s interface was partly inspired by the computational notebook paradigm. Just like how notebooks help break apart a script into smaller components for unit testing and quick iteration, our “agent notebook” design—featuring plan step “cells” and options for stepwise or continuous execution—help bring similar benefits to agentic workflows. We saw in both our lab study (Section 5) and field deployment (Section 7) that participants valued the ability to execute steps one at a time and interactively refine outputs before moving onto subsequent steps. We now discuss agent notebooks as a potentially generalizable design pattern and consider its applications to agentic systems beyond Cocoa.
@@ 8.2. The Agent Notebook Paradigm?
Consider “deep research” systems that assemble a detailed report on a particular topic upon searching for and synthesizing information from around the web (Google, 2025; OpenAI, 2025b; Perplexity, 2025). These systems may clarify user requests in a chat conversation (OpenAI, 2025b), or allow the user to prompt the agent to regenerate the plan (Google, 2025) before starting the research process. However, there is no room for user involvement thereafter. Agent notebooks can be particularly valuable here because they afford iteration and control over specific parts of planning and execution. For example, the user may not be satisfied with the sources the agent is using to inform its answers and wants to steer the agent to use information from a narrow set of websites for a particular plan step. This kind of interaction is expensive without agent notebooks because it requires re-running the entire research process. Agent notebooks allow the user to focus on improving outcomes of a particular step by editing that step and rerunning it or directly editing the agent’s outputs at that step, which can save significant time and cost.
While we envision agent notebooks to be a design pattern that is generalizable beyond Cocoa, we note that the specific affordances for co-planning and co-execution may vary depending on the environment in which human-agent collaboration happens. Cocoa’s environment is a document editor for scientific research, and contains affordances designed around that environment (e.g., representing plan steps as editable bulleted lists, paper curation and search in the sidebar). Translating these affordances to coding, for example, may need to draw upon different interaction metaphors. The plan can be represented as an executable spec file (e.g., a special Jupyter notebook with only markdown cell blocks that serve as step instructions), while co-execution can resemble live pair programming. We anticipate that making intentional design choices for these affordances by drawing from appropriate metaphors will be an important part of future work.
8.3. The Practical Significance of User Steps and Planning Ahead
@@
In Cocoa, users are presented with an interactive multi-step plan up front, which the agent adheres to and can modify while operating. Importantly, this plan also allows users to flexibly delegate a step to themselves or the agent. Both design decisions are a departure from many existing agentic AI systems, where the agent dynamically constructs a plan step-by-step as it generates outputs (Yao et al., 2022b) and seeks user input only when it automatically determines it should do so (OpenAI, 2025a; Peng et al., 2025; Shao et al., 2024). We discuss two practical reasons—cost and oversight—for why such a departure may be desirable for practical agents in the real world.
@@
AI agents can be prohibitively expensive to run (Kapoor et al., 2024; Yang et al., 2024). Because an agent’s actions often involve multiple LLM calls, costs can quickly accumulate, especially for more complex and long-running tasks. In response, academic projects have implemented cost-capping measures to alleviate this (Yang et al., 2024), and researchers have called for agent evaluations to be cost-controlled (Kapoor et al., 2024). In user-facing interactive systems, cost also includes time spent waiting for agents to come back with results. These cost considerations change user behavior. When participants made the decision to assign most or all steps to the agent in our lab studies, they viewed the agent’s actions as rather inconsequential and low-cost—they could run the agent and see what happens, and edit the output post hoc (Section 6.2). However, while this was the case in the lab, it may not be true in some real-world contexts where users and/or developers need to pay monetary and time costs for each agent step. User steps become an important cost-saving measure. Users can choose to complete a step that is not particularly burdensome or is too difficult for current AI agents to complete effectively. Efficiently leveraging human effort and expertise where appropriate via user steps can be a key step towards developing cost-aware agents that are usable and affordable in practice.
@@
The increased autonomy of AI agents also comes with increased risks, as AI harms become more difficult to anticipate, and accountability for AI actions becomes harder to trace (Chan et al., 2023, 2024). Cocoa’s interface not only provides more transparency into agent actions, but also provide some assurance about and control over the agent’s execution trajectory. By generating a plan ahead of execution that the agent adheres to, harm anticipation and mitigation become more tractable. Step-by-step plan generation used by existing agent frameworks, such as ReAct (Yao et al., 2022b), may also accomplish a similar goal by asking for user approval whenever a new step is generated. However, this can severely limit agent autonomy when desirable and may also disengage the user if their only form of supervision is repetitive rubber-stamping (Kapoor et al., 2024). Additionally, advanced planning enables assessments of task risk pre-execution. Plan steps determined to be of higher risk (e.g., requiring working with passwords) can be assigned to the user—even if the agent is capable of completing them—as a risk mitigation measure.